OpenStudy (anonymous):

PLEASE HELP! WILL GIVE A FAN and MEDAL!!!!!! Part 1: Create a scenario for an arithmetic sequence. For example, Jasmine practices the piano for ______ minutes on Monday. Every day she ___________ her practice time by _________. If she continues this pattern, how many minutes will she practice on the 7th day? Be sure to fill in the blanks with the words that will create an arithmetic sequence. Use your scenario to write the function for the 7th term in your sequence using sequence notation.

OpenStudy (igreen):

Arithmetic Sequences add values. @Chocolate1031

OpenStudy (igreen):

So you could say something like she practices for 10 minutes on Monday, she adds her practice time by 2.

OpenStudy (igreen):

Just keep adding 2 to 10 to get the 7th term.

OpenStudy (igreen):

10 is the first term. 10 + 2 = 12 12 is the second term. 12 + 2 = 14 14 is the third term. Etc..

OpenStudy (anonymous):

I know, I could have done that long time ago, though I need to write the arithmetic sequence notation. So far I know arithmetic sequence notation would be: \[a _{n}=a _{1}+d(n-1)\]

OpenStudy (anonymous):

\[a _{1}\]is the first term of the sequence \[a _{n}\]is the value of the nth term of the sequence \[n\] is the number of terms in the sequence

OpenStudy (igreen):

Yes, that's correct.

OpenStudy (anonymous):

I just don't know how to put it all in the format of the equation based on the problem

OpenStudy (anonymous):

are you still there?

OpenStudy (igreen):

First blank would be y-intercept Second blank would be type of slope like positive or negative Third blank would be value of slope

OpenStudy (anonymous):

Can you explain that in the equation format, please?

OpenStudy (igreen):

Okay, hold on.

OpenStudy (igreen):

I think this is the Sequence Notation: \(S_n=S(r)^{n+1}\)

OpenStudy (anonymous):

Are you sure? I thought it was \[S _{n}=\frac{ a _{1}(r ^{n})-a _{1} }{ r-1 }\]

OpenStudy (igreen):

Agh..hold on.

OpenStudy (igreen):

Oh wait..that's for Geometric Sequences..my bad.

OpenStudy (anonymous):

So what do I do now?

OpenStudy (igreen):

Yeah, this is it: \(a _{n}=a _{1}+d(n-1)\)

OpenStudy (igreen):

Where: a1 = the first term of the sequence an = the value of the nth term of the sequence n = the number of terms in the sequence r = the common ratio between the terms

OpenStudy (igreen):

So you have: Day 7 = 10 + 2(7-1)

OpenStudy (igreen):

That's the Sequence Notation.

OpenStudy (igreen):

Sorry it took me a while to find it.

OpenStudy (anonymous):

No problem, thank you so much! your like so much better than any teacher or lesson. :) Hold on let me give you a fan and a medal

OpenStudy (igreen):

Lol, I hate learning from lessons too, It's better to learn from people.

OpenStudy (anonymous):

ya I can't believe I am saying this, but I can't wait till school starts

OpenStudy (igreen):

Oh, and that Sequence notation is ONLY for: For example, Jasmine practices the piano for \(\color{red}{TEN}\) minutes on Monday. Every day she \(\color{red}{INCREASES}\) her practice time by \(\color{red}{TWO}\). If she continues this pattern, how many minutes will she practice on the 7th day? Be sure to fill in the blanks with the words that will create an arithmetic sequence. Use your scenario to write the function for the 7th term in your sequence using sequence notation.

OpenStudy (igreen):

I'm all-year round.

OpenStudy (igreen):

If you chose different values: Day 7 = 10 + 2(7-1) ^ = First Blank Day 7 = 10 + 2(7-1) ^ = Third Blank Everything else stays the same. Just make sure you say it increases(Second Blank).

OpenStudy (anonymous):

yes I know that because I read it in the lesson, but thanks for reminding me!

OpenStudy (igreen):

No problem. You're doing FLVS, right? I remember doing this..

OpenStudy (anonymous):

yes, I am taking algebra one honors

OpenStudy (igreen):

I finished that 2 or 3 months ago :D

OpenStudy (igreen):

Just tag me if you need any help.

OpenStudy (anonymous):

k thx for ur help.